Two trains are traveling toward eachother on adjacent tracks. ONe of the trains is 230 ft long and is moving at a speed of 50 m/h. The other train is 210 ft long and is moving at a rate of 60 m/h. Find the time interval between the moment the trains first meet until they completely pass each other. (Remember 1 mile= 5280feet}
Answer by mananth(16949)

Length of train I = 230 ft.
length of second train = 210 ft.
Length of both trains = 230 +210 = 440 ft.
This is the distance to be traveled so that they cross each other completely.
1 mile = 5280 ft
440 ft = 440/5280 = 0.083 miles
..
The trains are moving towards each other. add up speeds
50 + 60 = 110 mph
time = distance / speed
time = 44/528 * 1/110
Time = 1/1320 hours
= 1/1320 * 3600
= 2.73 seconds
They wil take 2.73 second to cross each other